OpenStudy (anonymous):

What is 1/0.1 ? I DONT WANT THE VALUE. just the pure meaning of it. like 2*3 means 2+2+2. AND 1/2 means a thing which is divided into two equal parts.

OpenStudy (anonymous):

ALSO PLZ ANY ONE CAN EXPLAIN ME WHAT 0/0 really means.

OpenStudy (ganpat):

It is a simple division... and similar to 1/2... as u know, 1/2 is dividing into half.. 1/ 0.1 = dividing into some x parts , each of 0.1 units... where, x = 1/0.1 = 10...

OpenStudy (anonymous):

YEAH NOW THAT IS WHERE I GOT PROBLEM. IN 1/2 it is divided into 2 parts But in 1/0.1 it is divide into 10 parts. shouldn't it also be 0.1 parts and size of each part is 10 which is more than the initial quantity.

OpenStudy (anonymous):

0/0 can be any number. Let's look at it from a different perspective. Let x be the number x = 0/0 0x = 0 ^ see what I did there?

OpenStudy (anonymous):

I also used to think the same way. But now I came to know that is wrong. @georgeblue22

OpenStudy (anonymous):

how is it wrong then?

OpenStudy (anonymous):

how can u cut nothing into nothing pieces and get something. DOES not make sense right?

OpenStudy (anonymous):

it does in this case. if you take a number and divide by 0, it is undefined. x = 1/0 0x = 1 ??? but the only exception is 0/0 sometimes maths is like that, hard to predict, but we must always look at it mathematically

Parth (parthkohli):

How can a number which is undefined be operated on like normal numbers?

Parth (parthkohli):

Undefined numbers cause all the problems! Let me prove that \(2 = 1\). Let's assume that \(a = b\). \(a^2 = ab\) \(a^2 - b^2 = ab - b^2\) Factor. \((a + b)(a - b) = b(a - b)\) Divide both sides by \(a - b\). \(a + b = b\) Since \(a = b\), we may substitute \(a\) with \(b\). \(b + b = b\) \(2b = b\) Divide both sides by 2. \(2 = 1\)

OpenStudy (anonymous):

^that's because b = 0

Parth (parthkohli):

The explanation is that \(a - b = 0\), and we can't divide both sides by \(0\).

Parth (parthkohli):

Remember that the number \(0\) either causes problems or isn't necessary in isolation.

OpenStudy (ganpat):

ha ha, ohkey , this is explanation for what i understand from your question... whatever term used in bottom, it says that u dividing ( cutting down ) the original piece in that number of parts... and what u get after solving is size of each part... eg: 1/2 = 0.5 Now 1/1 = 1... there is no cut.. everything on its place.. anything goes below that, the statement still remains same.. except what u dividing is size of each part and the answer u get is number of pieces.. eg: 1/0.5 = 2 Hope this answers ur question...

Parth (parthkohli):

• Don't divide by 0, as division by 0 results both sides to become infinite. • Don't multiply by 0, as both sides will become equal even if they are not! • Don't add/subtract by 0, as it doesn't help to do so(you'd just get the same equality even if you do so).

Parth (parthkohli):

Hmm, and the answer to the original question. \(0.1 \times \text{some number} = 1\) That's an oral question, because we just multiply 10 to shift the decimal one to the left. :)

OpenStudy (anonymous):

It means 1 is divided into 10 equal parts each of length 0.1..

OpenStudy (anonymous):